# Scalable multiphoton generation from cavity-synchronized single-photon sources

### v2

@@ V Inhomogeneity effects
Therefore, frequency inhomogeneties should also be controlled $\delta\omega_{q}^{j}\ll\Omega_{\rm cm}$, in order to have a high quality synchronization and demultiplexing. We discussed in Sec. III.1 that the qubit frequencies can be fine-tuned by sending specific DC currents on each antenna channel. With this we can make all qubits nearly resonant up to a tuning imprecision in the range $\Delta\omega_{q}/2\pi\lesssim 1$MHz DiCarlo et al. (2009); Barends et al. (2013); Pechal et al. (2014); Arrangoiz-Arriola et al. (2019). This means that a frequency disorder of order $\Delta\omega_{q}/\Omega_{\rm cm}\lesssim 2.5\%$ is achievable with state-of-the-art technology. In Figs. 9(c)-(d), we calculate $\langle\!\langle P_{N}\rangle\!\rangle_{\omega_{q}}$ and $\langle\!\langle D_{N}\rangle\!\rangle_{\omega_{q}}$ up to $N=4$ and for frequency disorder strengths $\Delta\omega_{q}/\Omega_{\rm cm}\lesssim 0\%$ (blue), $\Delta\omega_{q}/\Omega_{\rm cm}\lesssim 1\%$ (red), and $\Delta\omega_{q}/\Omega_{\rm cm}\lesssim 2.5\%$ (green). These results confirm that under these realistic disorder conditions the $N$-photon generation efficiency $\langle\!\langle P_{N}\rangle\!\rangle_{\omega_{q}}$ is minimally altered with respect to the homogeneous prediction [cf. blue data in Fig. 9(c)]. Moreover, we observe that the demultiplexing error $\langle\!\langle D_{N}\rangle\!\rangle_{\omega_{q}}$ reduces with higher frequency disorder $\Delta\omega_{q}/\Omega_{\rm cm}$ due to the larger independence of the photon emission processes. This occurs at the expense of reducing the generation efficiency $\langle\!\langle P_{N}\rangle\!\rangle_{\omega_{q}}$ so it is not a good limit for our purpose.
The synchronized multiphoton generation scheme is thus resilient to moderate disorder in all system paremeters and we expect high $N$-photon generation efficiencies in realistic circuit-QED implementations of the device.
VI Conclusions and Outlook
@@
In this work, we propose a scalable design for generating efficiently a large number of synchronized, independent and indistinguishable photons propagating over independent channels. The synchronization is provided by a strongly driven resonator in the bad cavity and weak-coupling limits. The resonator allows for a simultaneous and robust control of all emitters, which deposit photon into individual waveguides, at a high rate and with a high collection efficiency. Although our scheme can be implemented in cavity QED and nanophotonic platforms Welte et al. (2018); Casabone et al. (2015); Araneda et al. (2018); Reitz et al. (2013); Lodahl et al. (2015); Antón et al. (2019); Ellis et al. (2018); Goban et al. (2015), we have discussed a particularly efficient superconducting circuits implementation, with microwave resonators, superconducting qubits and microwave transmission lines Pechal et al. (2014); Zhou et al. (2020); Wang et al. (2020); Barends et al. (2013); Caldwell et al. (2018); McKay et al. (2019); Andersen et al. (2020); Dassonneville et al. (2020a).
@@
The only intrinsic limitation for the scalability of the synchronized multiphoton device are the cavity-mediated interactions and collective decay, which can create correlations between SPSs and the emitted photons. Nevertheless, for state-of-the-art circuit QED parameters, we show that these correlations are strongly suppressed on the timescale of photon emission, and that they only induce a nearly negligible quadratic demultiplexing error $D_{N}\approx\epsilon N(N-1)$ with $\epsilon\sim 10^{-8}-10^{-6}$. Remarkably, this allows for the synchronization of up to hundreds of nearly independent single-photons, even in the presence of decoherence and disorder.
@@
Given that, as we show, each SPS can achieve single-photon efficiency, purity, and indistinguishability above $99\%$, and the parallel operation in our device enables the efficient creation of large $N$-photon states. For instance, we predict a $30$-photon probability of $P_{30}\approx 0.72$ at a rate $C_{30}\approx 270$kHz and a $100$-photon probability of $P_{100}\approx 0.33$ at a rate $C_{100}\approx 130$kHz. This is seven orders of magnitude more efficient than the most sophisticated multiplexed SPSs up to $N=14$ Lenzini et al. (2017); Wang et al. (2017); Loredo et al. (2017); Wang et al. (2019a); Antón et al. (2019); Hummel et al. (2019). All these promising figures of merit can be further improved when implementing a more refined model for the SPSs such as three-level emitters Fischer et al. (2016), but this is independent from the efficient synchronization scheme we propose.
@@
Scalable and deterministic sources of multi-photon states will be a key resource for realizing large-scale quantum information processing such as quantum optical neural networks Steinbrecher et al. (2019) or fault-tolerant photonic quantum computation Takeda and Furusawa (2019). In the short-term, the implemenation of our scheme can already enable quantum advantage experiments with hundreds of microwave photons such as boson sampling Brod et al. (2019) or quantum metrology Motes et al. (2015); Ge et al. (2018). Moreover, the setup and ideas introduced in this work can be further extended to scale up the generation of correlated multi-photon states Meyer-Scott et al. (2019); Pfaff et al. (2017) with engineered entanglement patterns Besse et al. (2020). Many-body methods such as Matrix-product-states Huang et al. (2019); Orús (2014) could be exploited to study the multi-photon correlations of the propagating fields, and recent multi-photon probing methods Ramos and García-Ripoll (2017); Muñoz et al. (2018); Dassonneville et al. (2020b); Lescanne et al. (2020); Le Jeannic et al. (2021) could be use to characterize them in the lab.
@@ Acknowledgements
The authors thank S. Paraoanu, F. Luis, R. Martínez, R. Dassonneville, O. Buisson, Z. Wang, E. Torrontegui, M. Pino, A. González-Tudela, and D. Porras, for helpful discussions. Work in Madrid was supported by project PGC2018-094792-B-I00 (MCIU/AEI/FEDER, UE), CAM/FEDER project No. S2018/TCS-4342 (QUITEMAD-CM), and CSIC Quantum Technology Platform PT-001. T.R. further acknowledges funding from the EU Horizon 2020 program under the Marie Skłodowska-Curie grant agreement No. 798397. M.L. acknowledges support from the China Scholarship Council and National Natural Science Foundation of China under Grant No. 11475021.
@@ Appendix A Derivation of effective model for cavity synchronization
In this appendix we outline the derivation of the effective master equation (12) by performing a displacement transformation on the cavity mode [cf. Sec. A.1], and then an adiabatic elimination of the cavity fluctuations [cf. Sec. A.2].
A.1 Coherent displacement of driven cavity mode

### v3

@@ V Inhomogeneity effects
Therefore, frequency inhomogeneities should also be controlled $\delta\omega_{q}^{j}\ll\Omega_{\rm cm}$, to have a high-quality synchronization and demultiplexing. We discussed in Sec. III.1 that the qubit frequencies can be fine-tuned by sending specific DC currents on each antenna channel. With this we can make all qubits nearly resonant up to a tuning imprecision in the range $\Delta\omega_{q}/2\pi\lesssim 1$MHz DiCarlo et al. (2009); Barends et al. (2013); Pechal et al. (2014); Arrangoiz-Arriola et al. (2019). This means that a frequency disorder of order $\Delta\omega_{q}/\Omega_{\rm cm}\lesssim 2.5\%$ is achievable with state-of-the-art technology. In Figs. 9(c)-(d), we calculate $\langle\!\langle P_{N}\rangle\!\rangle_{\omega_{q}}$ and $\langle\!\langle D_{N}\rangle\!\rangle_{\omega_{q}}$ up to $N=4$ and for frequency disorder strengths $\Delta\omega_{q}/\Omega_{\rm cm}=0\%$ (blue), $\Delta\omega_{q}/\Omega_{\rm cm}=1\%$ (red), and $\Delta\omega_{q}/\Omega_{\rm cm}=2.5\%$ (green). These results confirm that under these realistic disorder conditions the $N$-photon generation efficiency $\langle\!\langle P_{N}\rangle\!\rangle_{\omega_{q}}$ is minimally altered with respect to the homogeneous prediction [cf. blue data in Fig. 9(c)]. Moreover, we observe that the demultiplexing error $\langle\!\langle D_{N}\rangle\!\rangle_{\omega_{q}}$ reduces with higher frequency disorder $\Delta\omega_{q}/\Omega_{\rm cm}$ due to the larger independence of the photon emission processes. This occurs at the expense of reducing the generation efficiency $\langle\!\langle P_{N}\rangle\!\rangle_{\omega_{q}}$ so it is not a good limit for our purpose.
We conclude that the synchronized multiphoton generation scheme is resilient to moderate disorder in all system parameters and that the required level of homogeneity is achievable in state-of-the-art circuit-QED implementations with parameter sets A and B.
VI Multi-mode nature of the cavity
So far we have assumed a mono-mode model for the cavity that synchronizes all emitters. Nevertheless, when increasing the length of the cavity to accommodate more qubits, the free spectral range of the cavity reduces and thus more normal cavity modes can participate in the dynamics Sundaresan et al. (2015). In this section, we consider these multi-mode effects and show that they can be described with the same effective model in Eqs. (13)-(18), but with a redefinition of the cavity-mediated parameters $\Omega_{\rm cm}$, $\delta_{\rm cm}$, $J_{\rm cm}$, and $\gamma_{\rm cm}$ [cf. Sec. VI.1]. We analyze multi-mode effects for various cavity lengths and configurations [cf. Secs. (VI.2)-(VI.4)] and show how to fine-tune the parameters to recover the same synchronization dynamics discussed previously with a mono-mode model. Moreover, we find that the participation of more cavity modes can even be beneficial for the multi-photon generation scheme as it substantially reduces the required driving strength on the cavity.
VI.1 Effective model including the multi-mode contributions
To account for the multi-mode nature of the cavity, we generalize the Hamiltonian and master equation in Eqs. (5)-(7), by considering that the qubits interact with many cavity modes $m$, each of them described by a frequency $\omega_{c}^{m}$, a decay $\kappa_{j}$ and a qubit-cavity coupling $g_{j,m}$ (see App. A.1 for more details). These quantities scale with mode number $m\in\mathbb{N}$ as Sundaresan et al. (2015),
$\omega_{c}^{m}$ $=m\omega_{c}^{1},$ (36)
$\kappa_{m}$ $=m\kappa_{1},$ (37)
$g_{j,m}$ $=\sqrt{m}g_{j,1},$ (38)
with $\omega_{c}^{1}$, $\kappa_{1}$, and $g_{j,1}$ the frequency, decay, and coupling corresponding to the fundamental mode $m=1$. Note that these fundamental parameters in turn depend on the geometry and length $L$ of the cavity, such as $\omega_{c}^{1}=\pi c/L$ for an open transmission line resonator, and thus the free spectral range scales as $\sim 1/L$.
It is possible to find parameters for which all cavity modes are in the bad-cavity limit, i.e. they are off-resonant and weakly coupled to the qubits, $g_{j,m}\ll\kappa_{m},|\Delta_{m}|$, so that their dynamics can be adiabatically eliminated. Here, $\Delta_{m}=\omega_{d}-\omega_{c}^{m}$ is the detuning between cavity modes and the external drive $\omega_{d}$, which is taken to be nearly resonant to the emitters $\omega_{d}\approx\omega_{j}^{q}$. As shown in Appendices A.2-A.3, the effective dynamics of the qubits under the above conditions is indeed described by the same effective Master equation and Hamiltonian in Eqs. (13)-(14), but with modified cavity-mediated quantities given by
$\Omega_{\rm cm}^{j}={}$ $\sum_{\omega_{c}^{m}\sim\omega_{d}}\frac{\Omega_{0}g_{j,m}}{\sqrt{(\kappa_{m}/2)^{2}+\Delta_{m}^{2}}},$ (39)
$\gamma^{jl}_{\rm cm}={}$ $\sum_{\omega_{c}^{m}\sim\omega_{d}}\frac{g_{j,m}g_{l,m}\kappa_{m}}{(\kappa_{m}/2)^{2}+\Delta_{m}^{2}},$ (40)
$J^{jl}_{\rm cm}={}$ $\sum_{\omega_{c}^{m}\sim\omega_{d}}\frac{g_{j,m}g_{l,m}\Delta_{m}}{(\kappa_{m}/2)^{2}+\Delta_{m}^{2}},$ (41)
and $\delta^{j}_{\rm cm}=J_{\rm cm}^{jj}$.
These expressions are the multi-mode generalization of Eqs. (15)-(18). They take into account the contribution of all cavity modes, with the largest effect coming from the modes closer to the qubit resonance $\omega_{d}\sim\omega_{j}^{q}$. Therefore, the effect of the multi-mode nature of the cavity is to modify the effective parameters of the dynamics. However, it is always possible to fine-tune the parameters to obtain the same effective physics discussed previously in the manuscript. In the following, we show this on three different multi-mode cavity configurations, namely an open transmission line resonator of moderate length [cf. Fig. 10(a)], a very long transmission line resonator [cf. Fig. 10(b)], and a ring cavity [cf. Fig. 10(c)]. For simplicity, we perform the analysis for homogeneous emitters in the remainder of the section.
Figure 10: Multi-mode effects in the cavity-mediated synchronization of multiple single-photon sources. We consider three different configurations that lead to the same effective dynamics discussed in this work: (a) an open transmission line resonator of moderate length L=9.17L=9.17mm (solid) and L=8.46L=8.46mm (dashed), (b) a long open transmission line resonator with length L=662L=662mm, and (c) a ring cavity with length L=74.6L=74.6mm. For each of these configurations (a)-(c), panels (d)-(f) show the mode frequency $\omega_{c}^{m}$, the coupling/decay ratio $g_{m}/\kappa_{m}$, and the mean steady-state occupation $|\alpha^{\rm ss}_{m}|^{2}$ as a function of the mode index $m$. The pink horizontal line in the frequency plots indicates the transition frequency $\omega_{q}/2\pi$ of the single-photon emitters on each case to see which cavity modes are closer to resonance.
VI.2 Transmission line resonator of moderate size
In this subsection, we show that multi-mode effects are weak in the synchronization setup for cavities of moderate size and when the qubits couple close but still off-resonant to the fundamental resonator mode $m=1$. For this, we consider a transmission line resonator of length L=9.17L=9.17mm [cf. Fig. 10(a)], so that the fundamental frequency reads $\omega_{c}^{1}/2\pi=7.2$ GHz. All other parameters correspond to set A of Table 1, namely $\Delta_{1}/2\pi=1400\textrm{ MHz}$, $g_{1}/2\pi=4\textrm{ MHz}$, $\kappa_{1}/2\pi=50\textrm{ MHz}$, and $\Omega_{0}/2\pi=14$ GHz. Fig. 10(d) displays the mode frequencies $\omega_{c}^{m}$, the coupling/decay ratio $g_{m}/\kappa_{m}$, and the average occupation $|\alpha^{\rm ss}_{m}|^{2}$ as a function of mode label $m$. The pink horizontal line indicates the resonance frequency of the emitters, $\omega_{q}/2\pi=8.6$ GHz, to compare it with the mode frequencies. We see that all modes satisfy the bad-cavity limit with $g_{m}/\kappa_{m}<10^{-1}$ and that the occupation is indeed concentrated in the fundamental mode $|\alpha^{\rm ss}_{1}|^{2}\sim 100$ as it is closest to resonance. All other modes have occupations more than two orders of magnitude lower and they only slightly modify the effective quantities. Furthermore, if we consider a slightly shorter cavity of length $L=8.46$ mm and fine-tune the system parameters as $\Omega_{0}/2\pi=10.4\textrm{ GHz}$, $\kappa_{1}/2\pi=50\textrm{ MHz}$, and $\Delta_{1}/2\pi=2000\textrm{ MHz}$, we recover the same effective quantities obtained from the mono-mode model in Sec. III.1 within a $0.1\%$ error, except for $J_{\rm cm}$ which is even reduced by $28\%$. The modifications of the mode structure and couplings are indicated by the dashed lines in Fig. 10(d), where one can see that steady-state occupations are less concentrated in mode $m=1$. The participation of more modes with similar weights allows the system to reduce cavity-mediated couplings as well as the required cavity drive.
VI.3 Long transmission line resonator
In this subsection, we show that the multi-mode structure of the cavity can be exploited to substantially reduce the driving strength required for parameters C and D from $\Omega_{0}/2\pi\sim 40$ GHz down to $\Omega_{0}/2\pi\sim 2.64$ GHz, while still inducing the same effective synchronization dynamics on the emitters. The idea is to consider a long resonator so that the free spectral range reduces and more modes can contribute to the effective parameters. In particular, we consider a long open transmission line resonator of length L=662L=662mm [cf. Fig. 10(b)], which can fabricated by bending the transmission line inside the chip Sundaresan et al. (2015); Chang et al. (2020). The length of this resonator is roughly $70$ times larger than in the previous subsection so the mode $m=50$ of frequency $\omega_{c}^{50}/2\pi=5$ GHz becomes the closest to resonance [cf. left panel of Fig. 10(e)]. Here, we consider qubits of frequency $\omega_{q}/2\pi\approx 4.98$ GHz, and all other parameters are chosen to induce the same effective dynamics as in the mono-mode model with parameters C, namely $g_{50}/2\pi=0.1\textrm{ MHz}$, $\kappa_{50}/2\pi=10.9\textrm{ MHz}$, $\Delta_{50}/2\pi=17.1\textrm{ MHz}$, and $\Omega_{0}/2\pi=2.64\textrm{ GHz}$. Note that the driving strength $\Omega_{0}$ is more than an order of magnitude lower than required in the modo-mode configuration [cf. Fig. 10(a)] This is because of the nearly $10$ off-resonant modes that play a non-negligible role in contributing to the effective parameters [cf. right panel of Fig. 10(e)], and in particular to the cavity-mediated driving $\Omega_{\rm cm}=\sum_{m}g_{m}|\alpha_{\rm ss}^{m}|^{2}\sim 2\pi\cdot 39$ MHz. Although we have discussed that having a strong resonator drive on order $\Omega_{0}/2\pi\sim 40$GHz, as required for parameter set C, is not harmful to the superconducting system (as most of this energy is reflected by the cavity), it is instructive to show that the multi-mode structure of the cavity does not introduce detrimental effects and that it may be even beneficial for reducing drive power if needed.
VI.4 Ring resonator
In Sec. V, we discussed that the multi-photon synchronizing scheme is resilient to disorder in the couplings $g_{j}$ as long as this is controlled with a relative precision on the order or below $\sim 2.5\%$. In this subsection, we consider a ring cavity configuration [cf. Fig. 10(c)], which is a very appealing alternative to improve the fabrication of homogeneous couplings $g_{j,m}\approx g_{m}$ due to its intrinsic circular symmetry. Prototypes of ring cavities Huang et al. (2021); Hazra et al. (2021) already demonstrate the generation of a very uniform electromagnetic field amplitude, leading to a coupling inhomogeneity below $1\%$. Note that for the same length $L$ of an open cavity, the ring cavity has a fundamental frequency twice as large $\omega_{c}^{1}=2\pi c/L$, due to its periodic boundary conditions. In particular, we consider a ring cavity of length $L=74.6\textrm{ mm}$, so that mode $m=3$ with $\omega_{c}^{3}/2\pi=5.31$GHz is the closest to the qubit resonance of $\omega_{q}/2\pi=6$ GHz [cf. left panel of Fig. 10(f)]. Other parameters are taken to induce an effective dynamics identical to parameters A in the mono-mode model, namely $\kappa_{3}/2\pi=15.6\textrm{ MHz}$, $g_{3}/2\pi=2.32\textrm{ MHz}$, $\Delta_{3}/2\pi=690\textrm{ MHz}$, and $\Omega_{0}/2\pi=3.7\textrm{ GHz}$. For these parameters, we see that the two modes $m=3,4$ are appreciably populated [cf. right panel of Fig. 10(f)]. As for the open cavity, increasing the length $L$ of the ring cavity also allows accommodating more qubits and reducing the driving strength $\Omega_{0}$ on the cavity, but the most important added feature of this symmetric configuration is the potential to reduce inhomogeneities in the fabrication of the multi-photon synchronizing device.
VII Conclusions and Outlook
@@
In this work, we propose a scalable design for generating efficiently a large number of synchronized, independent, and indistinguishable photons propagating over independent channels. The synchronization is provided by a strongly driven resonator in the bad cavity and weak-coupling limits. The resonator allows for a simultaneous and robust control of all emitters, which deposit photons into individual waveguides, at a high rate and with a high collection efficiency. Although our scheme can be implemented in cavity QED and nanophotonic platforms Welte et al. (2018); Casabone et al. (2015); Araneda et al. (2018); Reitz et al. (2013); Lodahl et al. (2015); Antón et al. (2019); Ellis et al. (2018), we have discussed an efficient circuit QED implementation, Pechal et al. (2014); Zhou et al. (2020); Wang et al. (2020); Barends et al. (2013); Caldwell et al. (2018), where the cavity synchronization mechanism is especially useful to reduce overhead with the number of independent controls Tamate et al. (2022).
@@
The only intrinsic limitations for the scalability of the synchronized multiphoton device are the cavity-mediated interactions and collective decay, which can create correlations between SPSs and the emitted photons. Nevertheless, we show that these correlations are strongly suppressed on the timescale of photon emission and that they only induce a nearly negligible quadratic demultiplexing error $D_{N}\approx\epsilon N(N-1)$ with $\epsilon\sim{\color[rgb]{0,0,0}10^{-4}-10^{-5}}$ for state-of-the-art circuit-QED parameters. Remarkably, this allows for the synchronization of up to hundreds of nearly independent single-photons, even in the presence of decoherence, disorder, and multi-mode cavity effects.
@@
Given that, as we show, each SPS can achieve single-photon efficiency, purity, and indistinguishability above $99\%$, and the parallel operation in our device enables the efficient creation of large $N$-photon states. For instance, we predict a $30$-photon probability of $P_{30}\sim 0.72$ at a rate $C_{30}{\color[rgb]{0,0,0}\sim 300}$kHz and a $100$-photon probability of $P_{100}\sim{\color[rgb]{0,0,0}0.16}$ at a rate $C_{100}\sim 200$kHz. This is seven orders of magnitude more efficient than the most sophisticated multiplexed SPSs up to $N=14$ single photons Lenzini et al. (2017); Wang et al. (2017); Loredo et al. (2017); Wang et al. (2019a); Antón et al. (2019); Hummel et al. (2019). All these promising figures of merit can be further improved when implementing a more refined model for the SPSs such as three-level emitters Fischer et al. (2016), but this is independent of the efficient synchronization scheme we propose.
@@
Scalable and deterministic sources of multi-photon states will be a key resource for realizing large-scale quantum information processing such as quantum optical neural networks Steinbrecher et al. (2019) or fault-tolerant photonic quantum computation Takeda and Furusawa (2019). In the short-term, the implementation of our scheme can already enable quantum advantage experiments with hundreds of microwave photons such as boson sampling Brod et al. (2019) or quantum metrology Motes et al. (2015); Ge et al. (2018). Moreover, the setup and ideas introduced in this work can be further extended to scale up the generation of correlated multi-photon states Meyer-Scott et al. (2022); Pfaff et al. (2017) with engineered entanglement patterns Besse et al. (2020). Many-body methods such as Matrix-product-states Huang et al. (2019); Orús (2014) could be exploited to study the multi-photon correlations of the propagating fields, and recent multi-photon probing methods Ramos and García-Ripoll (2017); Sánchez-Muñoz et al. (2018); Dassonneville et al. (2020); Lescanne et al. (2020); Le Jeannic et al. (2021) could be used to characterize them in the lab.
@@ Acknowledgements
The authors thank S. Paraoanu, F. Luis, R. Martínez, R. Dassonneville, O. Buisson, Z. Wang, E. Torrontegui, M. Pino, A. González-Tudela, and D. Porras, for helpful discussions. Work in Madrid was supported by project PGC2018-094792-B-I00 (MCIU/AEI/FEDER, UE), Proyecto Sinergico CAM 2020 Y2020/TCS-6545 (NanoQuCo-CM), and CSIC Quantum Technology Platform PT-001. T.R. further acknowledges funding from the EU Horizon 2020 program under the Marie Skłodowska-Curie grant agreement No. 798397 and the Ramón y Cajal program RYC2021-032473-I, financed by MCIN/AEI/10.13039/501100011033 and the European Union NextGenerationEU/PRTR. M.L. acknowledges support from the China Scholarship Council and the National Natural Science Foundation of China under Grant No. 11475021.
@@ Appendix A Derivation of effective model for cavity synchronization
In this appendix, we outline the derivation of the effective master equation (13) starting from general dynamics describing $N$ emitters coupled to a multi-mode cavity [cf. Sec. A.1]. In particular, we perform a displacement transformation on all the cavity modes [cf. Sec. A.2], and then an adiabatic elimination of the cavity fluctuations [cf. Sec. A.3] as shown in the following.
A.1 Open quantum dynamics of emitters coupled to a multi-mode cavity
The master equation and Hamiltonian describing $N$ two-level emitters coupled to a driven multi-mode cavity read,
$\dot{\rho}(t)={}$ $-i[H(t),\rho]+\sum_{m}\kappa_{m}{\cal{D}}[a_{m}]\rho+\sum_{j=1}^{N}\gamma_{j}{\cal{D}}[\sigma_{j}^{-}]\rho$
$+\sum_{j=1}^{N}\gamma_{\rm loss}^{j}{\cal{D}}[\sigma_{j}^{-}]\rho+\sum_{j=1}^{N}2\gamma_{\phi}^{j}{\cal{D}},[\sigma_{j}^{+}\sigma_{j}^{-}]\rho,$ (42)
$H(t)={}$ $\sum_{m=1}\omega_{c}^{m}a_{m}^{\dagger}a_{m}+2\Omega(t)\sum_{m=1}(a_{m}+a_{m}^{\dagger})\cos(\omega_{d}t)$
$+{}$ $\frac{1}{2}\sum_{j=1}^{N}\omega_{j}^{q}\sigma_{j}^{z}+\sum_{j=1}^{N}\sum_{m=1}g_{j,m}(a_{m}+a_{m}^{\dagger})(\sigma_{j}^{-}+\sigma_{j}^{+}).$ (43)
Here, the system parameters are the same as in the mono-mode case given in Eqs. (5)-(7), except that the cavity is generalized to have many modes $m$, each one described by ladder operators ($a_{m}$, $a_{m}^{\dagger}$), a frequency $\omega_{c}^{m}$, a decay $\kappa_{j}$ and a coupling to the qubits $g_{j,m}$.
A.2 Coherent displacement of driven cavity modes